A certain field is a rectangle with a perimeter of 966 feet. The length is 199 feet more that the width. Find the width and length of the rectangular field.
Answer by TimothyLamb(4379)

L = w + 199
2L + 2w = 966
---
2L + 2w = 966
4w = 568
---
w = 142
L = 341
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php